OpenStudy (anonymous):

Determine whether or not each of the following equations is dimensionally homogeneous. Show your work. a) F=ma b) F=m(V^2/R) c) F(t_2-t_1)=m(V_2-V_1) d) F=mV I have no idea how to solve this.

OpenStudy (anonymous):

what is dimension of F?

OpenStudy (anonymous):

? Idk. All I know is the F= force(N)

OpenStudy (anonymous):

Ma M = kg a = m/s^2 Dimmenssional Formula = [MLT^-2]

OpenStudy (anonymous):

F is the product of mass and acceleration. Acceleration is measured in meters per second squared, that is, Length per time squared. So, the dimensional formula of F is F=M * L/ T^2 or expressed as \[F=[MLT ^{-2}]\]

OpenStudy (anonymous):

similarly work out for each side, if in any case the dimensional formuae do not match, it is NON HOMOGENEOUS, else it is homogeneous

OpenStudy (anonymous):

I am so confused... can someone take it [SLOW] LOL.. please :/

OpenStudy (anonymous):

umm, okk, what r the units of m and a? unit of F is N which is equalt to kgm/s2

OpenStudy (anonymous):

m=kg a=m/s^2 Do I just write: F=ma N=(kg)(m/s^2) (kg)(m/s^2)=kg(m/s^2) which means they are equal.....? is that what I do?

OpenStudy (anonymous):

um yes, but in a different way lol unit of F is kgm/s2 so its dimensional formula wil be : [F] =[ M L T^-2 ]

OpenStudy (anonymous):

[F] =[ M L T^-2 ] wth is that^? lol can you explain that.

OpenStudy (anonymous):

similarly, dimensional formula for ma will be : [ma] =[ M ] [ L T^-2 ] = [ M L T^-2 ] both r same, so dimensionally homogeneous

OpenStudy (anonymous):

well, u replace kg by M, m by L and s by T thats all ans use square brackets lol

OpenStudy (anonymous):

why do you replace it? do you have to?

OpenStudy (anonymous):

yes, thats called DIMENSIONAL ANALYSIS lol

OpenStudy (anonymous):

can u do it n show me for the b part?

OpenStudy (anonymous):

I'll try....

OpenStudy (anonymous):

F=m(V^2/R) N=(kg)((m/s^2)/(m)) [F]=[M][.... I'm stuck there.

OpenStudy (anonymous):

You don't necessarily have to change units to M,L,T (mass, length, and time), but you do need to break everything down into fundamental units so you have a consistent basis for comparison.

OpenStudy (anonymous):

Well, I understand the units and comparing them. But, once I start changing it to M,L,T.. that confuses me!

OpenStudy (anonymous):

B isn't homogeneous is it?

OpenStudy (anonymous):

it is. well as cliff said u can use untis only if that how u r taught in school? no need to convert then jus check if both sides have same untis

OpenStudy (anonymous):

F=m(V^2/R) N=(kg)((m/s^2)/(m)) (kg)(m/s^2)=(kg)((m/s^2)/(m)) How is that equal on both sides? Or did I do something wrong?

OpenStudy (anonymous):

Check this part: V² = ( m/s ) ² = m²/s²

OpenStudy (anonymous):

but there is still a (m) left...? both (kg) cancel out and both (m/s)^2 cancel out.. there is a (m) left?

OpenStudy (anonymous):

(kg)(m/s²)=(kg)((m²/s²)/(m)) = kg * m/s²

OpenStudy (anonymous):

Does it look equal now?

OpenStudy (anonymous):

how did you go from =(kg)((m²/s²)/(m)) =kg * m/s² ? there is 3 (m) m^2 and m

OpenStudy (anonymous):

m²/ m = m

OpenStudy (anonymous):

oh you simplified!